# Searching for Unparticles with the Cosmic Microwave Background

### 13 Mar 2026 v1

@@ I Introduction
Due to the high-dimensionality of contemporary CMB data (with $N_{\rm pix}\mathcal{O}(10^{7})$ pixels in Planck), we require sophisticated techniques to robustly extract the small inflationary bispectrum signatures. The standard work-horse is the Komatsu-Spergel-Wandelt (KSW) approach, which invokes factorizability of the underlying inflationary bispectrum to reduce the computation time of a bispectrum estimator from $\mathcal{O}(N_{\rm pix}^{3})$ to $\mathcal{O}(N_{\rm pix}\log N_{\rm pix})$. Many models, such as the unparticle model introduced above, are not inherently factorizable, thus we must proceed by an alternate (and usually approximate) path, such as measuring the bispectrum in a set of Legendre modes, Fourier-modes, or bins [e.g., 22, 23, 11, 38, 57, 12, 58]. Here, we adopt a different approach, which is sketched in Fig. 2. utilizing the machine learning framework introduced in [41] to obtain a highly accurate approximation to the target bispectrum of interest using a set of neural network basis functions. Unlike the conventional equilateral and orthogonal templates, our models depend on an a priori unknown parameter, the conformal scaling dimension $\Delta$. To efficiently characterize the broad parameter space, we supplement the factorization approach with a PCA-based compression (paying close attention to possible degeneracies with single-field bispectra), and implement the result in the CMB bispectrum code PolySpec. Finally, we apply our results to the latest Planck temperature and polarization dataset, which (a) places strong constraints on the unparticle scenario (optionally marginalizing over self-interactions), and (b) serves as a proof-of-concept for the analysis of general strongly-coupled (and non-factorizable) inflationary models.
@@
Figure 2: Schematic outline of this work. Given a set of primordial bispectrum templates, computed analytically using bootstrap techniques, we use linear optimization to obtain a small set of basis templates that describe a wide span of models. Using the techniques of [41], these are expanded in factorizable form using machine learning techniques, allowing them to be integrated in KSW-type CMB estimators. Finally, we use the PolySpec package to place constraints on the underlying template amplitudes, $f_{\rm NL}^{\Delta}$. Though we focus on unparticle templates in this work, this outline could be applied to any suite of models.
@@ II.3 Phenomenology
In this work, we will additionally consider the alternative orthogonal shape proposed in [54]:
@@
integrating over a triangular domain. To roughly mimic the sensitivity of Planck, we add a factor of $(k_{1}+k_{2}+k_{3})^{-1}$ to convert from three- to two-dimensions (following [22], and a set of weighting functions $W(k)$ to restrict to the relevant Fourier-space scales (with support for $2\times 10^{-4}\leq k/\mathrm{Mpc}^{-1}\leq 2\times 10^{-1}$), following [41].
@@ III.1 Orthogonalization
We represent a given unparticle template $S_{\Delta}$ as a linear combination of the equilateral and orthogonal templates, plus a sum of residual components expanded in a carefully-chosen basis (which will be specified below):
@@ III.2 Factorization
where $\alpha_{n},\beta_{n},\gamma_{n}$ are arbitrary functions and $w_{n}$ are scalar weights. Whilst the simplest templates (including those of (9)) can be expressed in this form, most realistic models cannot. To proceed, one must rewrite the desired model in a separable form, for example, expanding it in a modal basis [e.g., 22]. In this work, we adopt the approach of [41], which uses machine learning tools to construct a low-dimensional set of neural network basis functions, $\{\alpha_{n},\beta_{n},\gamma_{n}\}$, and weights, $\{w_{n}\}$, tailored to the target bispectrum. This has been shown to accurately reproduce a variety of bispectra using just a handful of terms (typically $N_{\rm fact}\lesssim 5$), which is far fewer than for the modal approach [e.g., 59]. Here, we apply these techniques to each of the SVD basis functions individually: an alternative approach would be to construct a joint basis directly from the $S_{\Delta}$ models, though this is more expensive to optimize.
@@
$\mathcal{L}({\rm basis})$ $=$ $\left\langle{S_{\rm fact}[{\rm basis}]-S_{\rm SVD}|S_{\rm fact}[{\rm basis}]-S_{\rm SVD}}\right\rangle/\left\langle{S_{\rm SVD}|S_{\rm SVD}}\right\rangle$ (17)
where $\left\langle{\cdot|\cdot}\right\rangle$ is the inner product defined in (12), and ‘basis’ contains the set of $3N_{\rm fact}$ one-dimensional functions used in (16), which are parametrized as neural networks with a single hidden layer. We introduce a number of improvements compared to the former work. Firstly, instead of learning the weights $\{w_{n}\}$ dynamically, we fix them to the maximum-likelihood solution, which is obtained via
$\left\langle{M_{n}|S_{\rm SVD}}\right\rangle=\sum_{n^{\prime}=1}^{N_{\rm fact}}w_{n^{\prime}}\left\langle{M_{n}|M_{n^{\prime}}}\right\rangle,\qquad M_{n}(k_{1},k_{2},k_{3})\equiv\alpha_{n}(k_{1})\beta_{n}(k_{2})\gamma_{n}(k_{3})+\text{5 perms.}$ (18)
This leads to faster training and a more accurate model. Secondly, we add a number of penalties to the loss function of (17) to ensure that the $N_{\rm fact}$ terms are approximately orthogonal and to avoid a blow-up of their amplitudes (noting the scaling degeneracies in (16)). In full, we include three terms:
@@
for $\mathcal{G}_{nn^{\prime}}\equiv\left\langle{M_{n}|M_{n^{\prime}}}\right\rangle$, which suppress off-diagonal contributions and keep the overall amplitudes close to unity. Whilst this correction is certainly ad hoc, we have found it to substantially improve the training efficacy.
@@ III.3 Estimation
To facilitate marginalization of self-interaction shapes, we add $f_{\rm NL}^{\rm orth}$ and $f_{\rm NL}^{\rm orth^{*}}$ estimators to PolySpec, building on the pre-existing $f_{\rm NL}^{\rm loc}$ and $f_{\rm NL}^{\rm eq}$ implementations. The former is a simple adjustment of the equilateral estimator (as is clear from the similarities in the templates of (9)) and matches previous works [e.g., 3]. To our knowledge, the $f_{\rm NL}^{\rm orth^{*}}$ estimator has not been previously implemented in CMB analyses owing to its more complex shape (10) [cf. 54], though remains possible since $S_{\rm orth^{*}}$ can be expressed as a sum of monomials and thus factorized explicitly:
@@
Our dataset is the Planck PR4 temperature anisotropy and $E$-mode polarization measurements, as obtained from the npipe processing pipeline [4] (matching a number of recent non-Gaussianity studies [e.g., 47, 41, 40, 30]). To avoid foreground contamination, we apply the common component separation mask in both temperature and polarization (inpainting small holes linearly), and work with component-separated maps obtained via the sevem prescription [2]. Data are filtered in harmonic-space using a translation-invariant power spectrum comprising the best-fit theory spectrum, the observed beam and transfer functions, and the noise (itself obtained from half-mission maps), and we restrict to $\ell\in[100,2048]$. To compute errorbars, we employ $100$ FFP10/npipe simulations [4, 1], with a further $100$ used to compute the linear term in (20).
@@ V Discussion
Finally, the technologies developed in this work can be straightforwardly applied to other primordial models. These could include additional strongly-coupled scenarios, such as the gapped unparticle model of [29], the holographic model of [37], or the more standard weakly-coupled collider [e.g., 16, 6, 34]. While we have focused on de Sitter invariant systems in this work, this is also not a limitation: all parts of the pipeline extend trivially to inflationary models with strong breaking of scale-invariance. As shown in [49, 29], unparticle exchange also yields novel signatures in the inflationary trispectrum; whilst a full-shape analysis along the lines of this work is not yet feasible, a simplified analysis focusing on factorizable signatures may be possible following [45, 47, 56]. We leave these extensions to future work.
@@ Acknowledgements.
This project arose from discussions at the 41st Annual Colloquium of the Institut d’Astrophysique de Paris. The computations in this work were run at facilities supported by the Scientific Computing Core at the Flatiron Institute, a division of the Simons Foundation. OHEP was inspired by the IAP’s nauticall division. GLP and CY are supported by Scuola Normale and by INFN (IS GSS-Pi). The research of GLP and CY is moreover supported by the ERC (NOTIMEFORCOSMO, 101126304). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. GLP is further supported by the Italian Ministry of Universities and Research (MUR) under contract 20223ANFHR (PRIN2022).

### 26 Mar 2026 v2

@@ I Introduction
Due to the high-dimensionality of contemporary CMB data (with $N_{\rm pix}\mathcal{O}(10^{7})$ pixels in Planck), we require sophisticated techniques to robustly extract the small inflationary bispectrum signatures. The standard work-horse is the Komatsu-Spergel-Wandelt (KSW) approach, which invokes factorizability of the underlying inflationary bispectrum to reduce the computation time of a bispectrum estimator from $\mathcal{O}(N_{\rm pix}^{3})$ to $\mathcal{O}(N_{\rm pix}\log N_{\rm pix})$. Many models, such as the unparticle model introduced above, are not inherently factorizable, thus we must proceed by an alternate (and usually approximate) path, such as measuring the bispectrum in a set of Legendre modes, Fourier-modes, or bins [e.g., 22, 23, 11, 39, 58, 12, 59]. Here, we adopt a different approach, which is sketched in Fig. 2. utilizing the machine learning framework introduced in [42] to obtain a highly accurate approximation to the target bispectrum of interest using a set of neural network basis functions. Unlike the conventional equilateral and orthogonal templates, our models depend on an a priori unknown parameter, the conformal scaling dimension $\Delta$. To efficiently characterize the broad parameter space, we supplement the factorization approach with a Principal Component Analysis (PCA) compression (paying close attention to possible degeneracies with single-field bispectra), and implement the result in the CMB bispectrum code PolySpec. Finally, we apply our results to the latest Planck temperature and polarization dataset, which (a) places strong constraints on the unparticle scenario (optionally marginalizing over self-interactions), and (b) serves as a proof-of-concept for the analysis of general strongly-coupled (and non-factorizable) inflationary models.
@@
Figure 2: Schematic outline of this work. Given a set of primordial bispectrum templates, computed analytically using bootstrap techniques, we use linear optimization (a principal component analysis; PCA) to obtain a small set of basis templates that describe a wide span of models. Using the techniques of [42], these are expanded in factorizable form using machine learning techniques, allowing them to be integrated in KSW-type CMB estimators. Finally, we use the PolySpec package to place constraints on the underlying template amplitudes, $f_{\rm NL}^{\Delta}$. Though we focus on unparticle templates in this work, this outline could be applied to any suite of models.
@@ II.3 Phenomenology
In this work, we will additionally consider the alternative orthogonal shape proposed in [55] (which we label $\mathrm{orthogonal}^{*}$):
@@
integrating over a triangular domain. To roughly mimic the sensitivity of Planck, we add a factor of $(k_{1}+k_{2}+k_{3})^{-1}$ to convert from three- to two-dimensions (following [22]), and a set of weighting functions $W(k)$ to restrict to the relevant Fourier-space scales (with support for $2\times 10^{-4}\leq k/\mathrm{Mpc}^{-1}\leq 2\times 10^{-1}$), following [42].
@@ III.1 Orthogonalization
We represent a given unparticle template $S_{\Delta}$ as a linear combination of the equilateral and orthogonal templates, plus a sum of residual components expanded in a carefully-chosen basis (obtained via a singular value decomposition (SVD), as specified below):
@@ III.2 Factorization
where $\alpha_{n},\beta_{n},\gamma_{n}$ are arbitrary functions and $w_{n}$ are scalar weights. Whilst the simplest templates (including those of (9)) can be expressed in this form, most realistic models cannot. To proceed, one must rewrite the desired model in a separable form, for example, expanding it in a modal basis [e.g., 22]. In this work, we adopt the approach of [42], which uses machine learning tools to construct a low-dimensional set of neural network basis functions, $\{\alpha_{n},\beta_{n},\gamma_{n}\}$, and weights, $\{w_{n}\}$, tailored to the target bispectrum. This has been shown to accurately reproduce a variety of bispectra using just a handful of terms (typically $N_{\rm fact}\lesssim 5$), which is far fewer than for the modal approach [e.g., 60]. Here, we apply these techniques to each of the SVD basis functions individually: an alternative approach would be to construct a joint basis directly from the $S_{\Delta}$ models, though this is more expensive to optimize.
@@
$\mathcal{L}({\rm basis})$ $=$ $\frac{\left\langle{S_{\rm fact}[{\rm basis}]-S_{\rm SVD}|S_{\rm fact}[{\rm basis}]-S_{\rm SVD}}\right\rangle}{\left\langle{S_{\rm SVD}|S_{\rm SVD}}\right\rangle}$ (17)
where $\left\langle{\cdot|\cdot}\right\rangle$ is the inner product defined in (12), and ‘basis’ contains the set of $3N_{\rm fact}$ one-dimensional functions used in (16), which are parametrized as neural networks with a single hidden layer. We introduce a number of improvements compared to the former work. Firstly, instead of learning the weights $\{w_{n}\}$ dynamically alongside $\alpha_{n},\beta_{n},\gamma_{n}$, we fix them to the maximum-likelihood solution (obtained by setting $\delta\mathcal{L}/\delta w_{n}=0$). The optimal weights are thus
$\left\langle{M_{n}|S_{\rm SVD}}\right\rangle=\sum_{n^{\prime}=1}^{N_{\rm fact}}w^{\rm opt}_{n^{\prime}}\left\langle{M_{n}|M_{n^{\prime}}}\right\rangle,\qquad M_{n}(k_{1},k_{2},k_{3})\equiv\alpha_{n}(k_{1})\beta_{n}(k_{2})\gamma_{n}(k_{3})+\text{5 perms.}$ (18)
This leads to faster training and a more accurate model. Secondly, we add a number of penalties to the loss function of (17):
@@
for $\mathcal{G}_{nn^{\prime}}\equiv\left\langle{M_{n}|M_{n^{\prime}}}\right\rangle$. Respectively, these ensure that the overall amplitudes of the $N_{\rm fact}$ basis functions are close to unity (noting that scaling degeneracy with $w_{n}$), that the basis is approximately orthogonal (with $\mathcal{G}_{nn^{\prime}}\to 0$ for $n\neq n^{\prime}$), and that degenerate modes are avoided. Whilst this correction is certainly ad hoc, we have found it to substantially improve the training efficacy.
@@ III.3 Estimation
To facilitate marginalization of self-interaction shapes, we add $f_{\rm NL}^{\rm orth}$ and $f_{\rm NL}^{\rm orth^{*}}$ estimators to PolySpec, building on the pre-existing $f_{\rm NL}^{\rm loc}$ and $f_{\rm NL}^{\rm eq}$ implementations. The former is a simple adjustment of the equilateral estimator (as is clear from the similarities in the templates of (9)) and matches previous works [e.g., 3]. To our knowledge, the $f_{\rm NL}^{\rm orth^{*}}$ estimator has not been previously implemented in CMB analyses owing to its more complex shape (10) [cf. 55], though remains possible since $S_{\rm orth^{*}}$ can be expressed as a sum of monomials and thus factorized explicitly:
@@
Our dataset is the Planck PR4 temperature anisotropy and $E$-mode polarization measurements, as obtained from the npipe processing pipeline [4] (matching a number of recent non-Gaussianity studies [e.g., 48, 42, 41, 30]). To avoid foreground contamination, we apply the common component separation mask in both temperature and polarization (inpainting small holes linearly), and work with component-separated maps obtained via the sevem prescription [2]. Data are filtered in harmonic-space using a translation-invariant power spectrum comprising the best-fit theory spectrum, the observed beam and transfer functions, and the noise (itself obtained from half-mission maps), and we restrict to $\ell\in[100,2048]$. To compute errorbars, we employ $100$ FFP10/npipe simulations [4, 1], with a further $100$ used to compute the linear term in (20).
@@ V Discussion
Finally, the technologies developed in this work can be straightforwardly applied to other primordial models. These could include additional strongly-coupled scenarios, such as the gapped unparticle model of [29], the five-dimensional model of [34], the holographic model of [38], or the more standard weakly-coupled collider [e.g., 16, 6, 35]. While we have focused on de Sitter invariant systems in this work, this is also not a limitation: all parts of the pipeline extend trivially to inflationary models with strong breaking of scale-invariance. As shown in [50, 29], unparticle exchange also yields novel signatures in the inflationary trispectrum; whilst a full-shape analysis along the lines of this work is not yet feasible, a simplified analysis focusing on factorizable signatures may be possible following [46, 48, 57]. We leave these extensions to future work.
@@ Acknowledgements.
This project arose from discussions at the 41st Annual Colloquium of the Institut d’Astrophysique de Paris. We thank Salvatore Sirletti for comments on the draft. The computations in this work were run at facilities supported by the Scientific Computing Core at the Flatiron Institute, a division of the Simons Foundation. OHEP was inspired by the IAP’s nauticall division. GLP and CY are supported by Scuola Normale and by INFN (IS GSS-Pi). The research of GLP and CY is moreover supported by the ERC (NOTIMEFORCOSMO, 101126304). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. GLP is further supported by the Italian Ministry of Universities and Research (MUR) under contract 20223ANFHR (PRIN2022).